# Holistic Explainable AI (H-XAI): Extending Transparency Beyond Developers in AI-Driven Decision Making

### v1

@@
Abstract
Current eXplainable AI (XAI) methods largely serve developers, often focusing on justifying model outputs rather than supporting diverse stakeholder needs. A recent shift toward Evaluative AI reframes explanation as a tool for hypothesis testing, but still focuses primarily on operational organizations. We introduce Holistic-XAI (H-XAI), a unified framework that integrates causal rating methods with traditional XAI methods to support explanation as an interactive, multi-method process. H-XAI allows stakeholders to ask a series of questions, test hypotheses, and compare model behavior against automatically constructed random and biased baselines. It combines instance-level and global explanations, adapting to each stakeholder’s goals, whether understanding individual decisions, assessing group-level bias, or evaluating robustness under perturbations. We demonstrate the generality of our approach through two case studies spanning six scenarios: binary credit risk classification and financial time-series forecasting. H-XAI fills critical gaps left by existing XAI methods by combining causal ratings and post-hoc explanations to answer stakeholder-specific questions at both the individual decision level and the overall model level.
1 Introduction
@@
Explainable AI (XAI) has long promised to make AI models more transparent and trustworthy. But despite the variety of methods available, most XAI tools are still designed for developers or model builders, not for the full range of stakeholders who interact with or are affected by AI systems (Mueller et al. 2019; Bhatt et al. 2020; Deshpande and Sharp 2022). Figure 1(b) illustrates three such key stakeholder groups: individual users (e.g., customers), operational organizations (e.g., data scientists), and regulatory bodies (e.g., policy-makers).
Consider the task of credit risk assessment. An applicant might reasonably ask, “Would my loan have been approved if I had requested a smaller amount?” A data scientist may wonder, “Which applicant features drive approval decisions on average?” A regulator might want to assess whether a model systematically favors certain demographic groups. These questions are all valid, but they differ in scope, require different types of explanations, and cannot be answered by a single method. Yet, current XAI methods are still geared toward justifying predictions rather than helping stakeholders explore models interactively. They typically do not support toggling between hypotheses (Hoffman et al. 2023b), testing counterfactuals (Miller 2023), or surfacing model limitations (Suresh et al. 2021). Most generate one-shot outputs, assuming explanation is a destination, not a process. But as any user who’s puzzled over a model’s behavior can attest, explanation is rarely a one-off affair (Hoffman et al. 2023a). Past work on black-box model rating has taken some steps toward user-centered analysis. (Srivastava and Rossi 2018) proposed a two-step method to rate AI services for bias, along with a tool for exploring trade-offs between accuracy and gender fairness in translation systems (Bernagozzi et al. 2021b). This approach, designed for settings where training data is unavailable, has since been applied to more complex systems like chatbots (Srivastava et al. 2020) and search engines (Tian et al. 2023). These efforts help communicate how models behave in the wild, but they still stop short of providing explanations aligned with stakeholder questions.
Causal fairness methods go further. They attempt to isolate the influence of protected attributes on the model’s predictions (Carey and Wu 2022). This aligns well with human reasoning: people often want to know whether changing a protected attribute, while holding everything else constant, would change the model’s decision. Building on this foundation, (Lakkaraju 2022; Srivastava et al. 2024) introduced a causal rating framework to evaluate black-box AI systems across domains, from sentiment analysis (Lakkaraju, Srivastava, and Valtorta 2024) to time-series forecasting (Lakkaraju et al. 2024, 2025). This approach quantifies both statistical and confounding bias and examines how input perturbations affect outputs. But while it helps isolate causal effects, it falls short when it comes to answering instance-level questions that matter in practice.
@@
In this paper, we propose a unified framework, Holistic-XAI (H-XAI), that integrates traditional XAI techniques with rating-driven explanations (RDE). RDE is well-suited for questions about model robustness, bias, and the effects of input perturbations, especially when those questions require comparisons across different models. In contrast, traditional methods such as SHAP, PDPs, and counterfactuals are more effective for explaining individual predictions or identifying which features drive decisions. H-XAI combines both, so stakeholders can move between overall model behavior and instance-level explanations, depending on the questions they care about. To ground our proposal, we present two case studies spanning six scenarios: credit risk classification using the German Credit dataset and financial time-series forecasting with stock prices. In both cases, we show how different types of questions posed by applicants, regulators, or data scientists can be addressed using appropriate combinations of methods.
Our contributions are as follows: 1. We introduce Holistic-XAI (H-XAI), a unified formulation that combines black-box ratings with post-hoc XAI techniques to support both individual- and system-level explanation needs (Figure 1(c)); 2. We propose a rating-driven explanation (RDE) workflow that allows hypothesis testing using causal metrics, and supports comparative model evaluation using automatically generated random and biased baselines (Figure 2); 3. We demonstrate how H-XAI supports a range of stakeholder needs through two representative case studies: binary classification and time-series forecasting, and summarize method selection strategies across questions and contexts (Section 4 and Appendix D).
@@
Figure 1: (a) Generalized Causal Diagram; (b) Stakeholder types in the context of Blackbox AI; (c) H-XAI Workflow: End-user receives output from a blackbox AI. Stakeholders may seek related explanations. “Method” shows type (.) and instances from case studies.
2 Background and Related Work
@@
In this section, we describe how we categorized stakeholders by drawing on prior work, and how we selected a subset of these stakeholder types to represent in our case studies in Section 4. We also explain how the questions we chose reflect the explanation needs identified in the literature, and conclude by situating our work within relevant prior work. An extended discussion of related work, including a dedicated section on explainability for time-series models, is available in the appendix (Section A).
2.1 Background: Stakeholders and Their Explanation Needs
@@
We follow the stakeholder taxonomy from (Deshpande and Sharp 2022), which identifies three main groups (Figure 1(b)):
@@
To structure their explanation needs, we adapt the XAI question taxonomy from (Liao et al. 2021), which organizes nine categories of explanatory questions from interviews with industry practitioners. Though originally domain-agnostic, many of their core questions, such as “Why was this prediction made?”, “What would change the outcome?”, or “Which features matter most?”, align closely with our stakeholder scenarios in credit scoring and forecasting. We pair each question type with a fitting method: SHAP for feature attribution, counterfactuals for what-if queries, and RDE for behavior under perturbations or across subgroups. A full mapping is provided in the appendix (Table D.2).
2.2 Explainability in AI
@@
Explainable AI (XAI) methods aim to clarify models behavior by describing predictions, decision patterns, and potential risks (Arrieta et al. 2020). Toolkits like AIX360 (Arya et al. 2019) offer diverse techniques to support this goal. However, most explanations remain static, built to justify outcomes rather than support interaction or hypothesis testing. Prior work argues that explanation is not a fixed output but a process of exploration (Mueller et al. 2019; Hoffman et al. 2023a). Users often need to ask follow-up questions or test alternate scenarios to build meaningful understanding (Abdul et al. 2018; Chromik and Butz 2021). Existing methods rarely support this kind of iterative reasoning or sensitivity analysis. In H-XAI, we use post-hoc methods like SHAP and PDPs to answer questions about feature influence on predictions, both locally and globally. They, however, are limited when users want to test hypotheses, such as whether a model’s decision would change under a counterfactual scenario, or whether outcomes differ systematically across groups. For such cases, we use RDE that estimate perturbation effects, account for confounding, and provide model comparisons.
2.3 Rating of AI Models
@@
Prior work has introduced rating methods to assess AI models from a third-party perspective, often without access to training data. (Srivastava and Rossi 2020) proposed a framework to quantify bias, particularly gender bias, in machine translation systems (Srivastava and Rossi 2018), using visualizations to communicate results (Bernagozzi et al. 2021b, a). However, these approaches lacked a causal foundation. Recent work introduced causal rating methods to isolate the effects of protected attributes on model outcomes. This was first applied to sentiment analysis systems (Lakkaraju, Srivastava, and Valtorta 2024) and then extended to composite NLP tasks (Lakkaraju et al. 2023) and time-series forecasting models (Lakkaraju et al. 2024, 2025). These methods go beyond correlation by estimating treatment effects and adjusting for confounding. In this paper, we build on that foundation but shift the role of ratings: rather than treating them solely as evaluation metrics, we use them to explain model behavior. By integrating ratings with traditional XAI methods, we allow stakeholders to compare models, test hypotheses, and interpret robustness and bias in a more structured way.
2.4 Understanding Stakeholder Needs in XAI
@@
Understanding stakeholder needs is critical for building explanations that are useful, not just technically accurate (Bhatt et al. 2020). For example, (Bundas et al. 2023) highlights NIST’s emphasis on hierarchical vocabularies: high-level concepts should be broadly accessible, while lower levels serve expert users. Studies show stakeholders prefer interactive explanations. Interviews in (Hoffman et al. 2023b) show a strong preference for tools that allow input manipulation and scenario exploration. RDE supports this by allowing users to perturb inputs and observe causal effects on model behavior (Figure D.3). While current XAI methods focus on local explanations, users often want a broader understanding of model behavior (Mueller et al. 2019). (Suresh et al. 2021) proposes a three-level framework for interpretability needs, including the need to understand model limitations. RDE addresses this by showing how behavior changes under different conditions, supporting users with a structured robustness evaluation. In this paper, we show how H-XAI brings these ideas together, supporting interactive, stakeholder-aligned explanations through a combination of rating and post-hoc XAI methods.
3 Approach
Figure 2: RDE Workflow.
@@ 3 Approach
3.1 Rating-Driven Explanation (RDE)
@@
This diagram allows us to define target causal estimands such as $E[O\mid do(T)]$ (i.e., the deconfounded distribution after adjustment), which represents the expected model outcome under an intervention on $T$. In practice, causal models are often specified by experts, derived from controlled experiments, or inferred using causal discovery algorithms (Gong et al. 2024; Varando et al. 2024), which themselves rely on statistical associations in the observed data. We begin with a hypothesized causal model that assumes protected attributes act as confounders and explore different scenarios within this structure. Even when based on assumptions, such models provide a structured framework for reasoning under interventions, something that correlational methods often fail to do, especially in the presence of confounding (as illustrated by Simpson’s paradox (Pearl 2022)). For instance, in Section 4.1, we consider the possibility that a loan applicant’s status influences the credit amount requested, an entirely plausible scenario in practice. The outcome $O$ may be either the model prediction $\hat{Y}$ or a derived quantity g(Y^)g($\hat${Y}). For example, in time-series settings, we define $O=|\hat{Y}-Y|$ as the residual error. This lets us ask: “Does a certain input perturbation systematically worsen performance for a particular group?”.
Rating Metrics
@@
While RDE builds on prior rating methods, it repositions their function. Earlier works primarily used ratings to compare model performance across input conditions. RDE instead uses them as part of a structured explanation workflow: formulating a question (e.g., “Does treatment T affect outcome O?”), selecting a baseline, applying a metric, and interpreting the result as a potential explanation. This gives ratings an explanatory role, especially in helping stakeholders reason about model bias, instability, or robustness under different conditions. We use three metrics, each designed to answer a specific kind of question; their complete definitions are provided in Appendix B.
1. Weighted Rejection Score (WRS): WRS quantifies statistical differences in outcomes across protected groups using Student’s $t$-test at multiple confidence levels (e.g., 95%, 75%, 60%). A weighted count of rejections across these thresholds yields the final score. WRS is especially useful in settings where the treatment is uniform (e.g., no variation in intervention across groups), such as in our forecasting use case (Section 4.2).
2. Average Treatment Effect (ATE): ATE estimates the causal effect of a treatment on model outcome, averaged over the population. We use propensity score matching (PSM) for deconfounding to improve estimation where needed (Rosenbaum and Rubin 1983).
3. Deconfounded Impact Estimation (DIE %): DIE % quantifies the degree of confounding by measuring the percentage difference between the unadjusted ATE and the ATE estimated after adjusting for the protected attribute (deconfounding).
These metrics serve different explanatory roles: WRS answers “Are outcomes distributed differently across protected attribute groups?”; ATE: answers “Does a treatment cause an outcome change?”; DIE %: answers “Is the observed effect real or driven by confounding?”
RDE Workflow

### v2

@@
Abstract.
As AI systems increasingly mediate decisions in domains such as credit scoring and financial forecasting, their lack of transparency and bias raise critical concerns for fairness and public trust. Existing explainable AI (XAI) approaches largely serve developers, focusing on model justification rather than the needs of affected users or regulators. We introduce Holistic eXplainable AI (H-XAI), a framework that integrates causality-based rating methods with post-hoc explanation techniques to support transparent, stakeholder-aligned evaluation of AI systems deployed in online decision contexts. H-XAI treats explanation as an interactive, hypothesis-driven process, allowing users, auditors, and organizations to ask questions, test hypotheses, and compare model behavior against automatically generated random and biased baselines. By combining global and instance-level explanations, H-XAI helps communicate model bias and instability that shape everyday digital decisions. Through case studies in credit risk assessment and stock price prediction, we show how H-XAI extends explainability beyond developers toward responsible and inclusive AI practices that strengthen accountability in sociotechnical systems.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Explainable AI (XAI) has long promised to help people understand how AI systems behave. Yet despite the variety of methods available, most existing tools are still designed for developers or model builders, not for the wider range of stakeholders who interact with or are affected by AI systems (Mueller et al., 2019; Bhatt et al., 2020; Deshpande and Sharp, 2022). Many such systems now appear in settings where decisions are delivered through online interfaces or web-based services, which further broadens who may need to interpret a system’s behavior. Figure 1(b) illustrates three key stakeholder groups: individual users (e.g., applicants using online decision tools), operational organizations (e.g., data science teams), and regulatory bodies.
Consider the task of credit risk assessment. An applicant might ask, “Would my loan have been approved if I had requested a smaller amount?” A data scientist may wonder, “Which applicant features drive approval decisions on average?” A regulator might want to know whether a model systematically favors certain demographic groups. These questions differ in scope and cannot be handled by a single explanation method. Current XAI tools typically focus on justifying a model’s output rather than helping stakeholders examine it from different angles. They rarely support moving between alternative hypotheses (Hoffman et al., 2023b), testing counterfactuals (Miller, 2023), or identifying when a model behaves unreliably (Suresh et al., 2021). They also tend to produce a fixed explanation, even though practical use often requires several rounds of inspection (Hoffman et al., 2023a).
Past work on black-box model rating offers one way to study model behavior without access to training data. (Srivastava and Rossi, 2018) introduced a two-step procedure to quantify bias in AI services, followed by visualization tools for exploring accuracy-fairness trade-offs (Bernagozzi et al., 2021b). These ideas have since been applied to chatbots (Srivastava et al., 2020) and search engines (Tian et al., 2023). While such ratings communicate how a system behaves “in the wild,” they do not answer the questions stakeholders often ask about why a particular decision changed or how sensitive the model is to certain inputs.
Causal fairness methods go a step further by isolating the influence of protected attributes on predictions (Carey and Wu, 2022). This mirrors the kinds of “what if this were different?” questions people tend to ask. Building on this line of work, (Lakkaraju, 2022; Srivastava et al., 2024) proposed a causal rating framework applied across domains ranging from sentiment analysis (Lakkaraju et al., 2024b) to time-series forecasting (Lakkaraju et al., 2024a, 2025). These methods quantify statistical and confounding bias and evaluate model stability under perturbations. However, they do not provide the instance-level explanations often needed in operational contexts.
@@
In this paper, we propose a framework, Holistic XAI (H-XAI), that brings together traditional XAI techniques and rating-driven explanations (RDE). RDE is helpful for questions involving model instability, bias, and comparisons across models; traditional methods such as SHAP, PDPs, and counterfactuals address local explanations and feature-level reasoning. H-XAI lets stakeholders move between these viewpoints depending on the question at hand. To illustrate how this works, we present two case studies covering six concrete scenarios: credit risk classification using the German Credit dataset and financial time-series forecasting with stock prices. Across these scenarios, we show how the combination of methods supports the kinds of questions raised by applicants, regulators, and data scientists.
Our contributions are as follows:
  1. (1)
    We introduce Holistic XAI (H-XAI), a formulation combining black-box ratings with post-hoc XAI techniques for both instance- and system-level analysis (Figure 1(c));
  2. (2)
    We develop a rating-driven explanation (RDE) workflow that supports hypothesis testing with causal metrics and comparative evaluation using automatically generated random and biased baselines (Figure 2);
  3. (3)
    We demonstrate how H-XAI addresses different stakeholder questions across binary classification and time-series forecasting (Section 4), and discuss practical method-selection strategies (Table 1).
(a) Causal Graph
(b) AI Stakeholders
(c) H-XAI Workflow
@@
Figure 1. (a) Generalized Causal Graph; (b) Stakeholder types in the context of Blackbox AI; (c) H-XAI Workflow: End-user receives output from a blackbox AI. Stakeholders may seek related explanations. “Method” shows type (.) and instances from case studies.
2. Background and Related Work
@@
In this section, we describe how we categorized stakeholders by drawing on prior work, and how we selected a subset of these stakeholder types to represent in our case studies in Section 4. We also explain how the questions we chose reflect the explanation needs identified in the literature, and conclude by situating our work within relevant prior work.
2.1. Background: Stakeholders and Their Explanation Needs
@@
We follow the stakeholder taxonomy from (Deshpande and Sharp, 2022), which identifies three main groups (Figure 1(b)):
@@
To structure their explanation needs, we adapt the XAI question taxonomy from (Liao et al., 2021), which organizes nine categories of explanatory questions from interviews with industry practitioners. Though originally domain-agnostic, many of their core questions, such as “Why was this prediction made?”, “What would change the outcome?”, or “Which features matter most?”, align closely with our stakeholder scenarios in credit scoring and forecasting. We pair each question type with a fitting method: SHAP for feature attribution, counterfactuals for what-if queries, and RDE for behavior under perturbations or across subgroups. A full mapping is provided in the appendix (Table A.2).
2.2. Explainability in AI
@@
Explainable AI (XAI) methods aim to clarify models behavior by describing predictions, decision patterns, and potential risks (Arrieta et al., 2020). Toolkits like AIX360 (Arya et al., 2019) offer diverse techniques to support this goal. However, most explanations remain static, built to justify outcomes rather than support interaction or hypothesis testing. Prior work argues that explanation is not a fixed output but a process of exploration (Mueller et al., 2019; Hoffman et al., 2023a). Users often need to ask follow-up questions or test alternate scenarios to build meaningful understanding (Abdul et al., 2018; Chromik and Butz, 2021). Existing methods rarely support this kind of iterative reasoning or sensitivity analysis. In H-XAI, we use post-hoc methods like SHAP and PDPs to answer questions about feature influence on predictions, both locally and globally. They, however, are limited when users want to test hypotheses, such as whether a model’s decision would change under a counterfactual scenario, or whether outcomes differ systematically across groups. For such cases, we use RDE that estimate perturbation effects, account for confounding, and provide model comparisons.
2.3. Rating of AI Models
@@
Prior work has introduced rating methods to assess AI models from a third-party perspective, often without access to training data. (Srivastava and Rossi, 2020) proposed a framework to quantify bias, particularly gender bias, in machine translation systems (Srivastava and Rossi, 2018), using visualizations to communicate results (Bernagozzi et al., 2021b, a). However, these approaches lacked a causal foundation. Recent work introduced causal rating methods to isolate the effects of protected attributes on model outcomes. This was first applied to sentiment analysis systems (Lakkaraju et al., 2024b) and then extended to composite NLP tasks (Lakkaraju et al., 2023) and time-series forecasting models (Lakkaraju et al., 2024a, 2025). These methods go beyond correlation by estimating treatment effects and adjusting for confounding. In this paper, we build on that foundation but shift the role of ratings: rather than treating them solely as evaluation metrics, we use them to explain model behavior. By integrating ratings with traditional XAI methods, we allow stakeholders to compare models, test hypotheses, and interpret robustness and bias in a more structured way.
2.4. Understanding Stakeholder Needs in XAI
@@
Understanding stakeholder needs is critical for building explanations that are useful, not just technically accurate (Bhatt et al., 2020). For example, (Bundas et al., 2023) highlights NIST’s emphasis on hierarchical vocabularies: high-level concepts should be broadly accessible, while lower levels serve expert users. Studies show stakeholders prefer interactive explanations. Interviews in (Hoffman et al., 2023b) show a strong preference for tools that allow input manipulation and scenario exploration. RDE supports this by allowing users to perturb inputs and observe causal effects on model behavior (Figure 12). While current XAI methods focus on local explanations, users often want a broader understanding of model behavior (Mueller et al., 2019). (Suresh et al., 2021) proposes a three-level framework for interpretability needs, including the need to understand model limitations. RDE addresses this by showing how behavior changes under different conditions, supporting users with a structured robustness evaluation. In this paper, we show how H-XAI brings these ideas together, supporting interactive, stakeholder-aligned explanations through a combination of rating and post-hoc XAI methods.
3. Approach
Figure 2. RDE Workflow.
@@ 3. Approach
3.1. Rating-Driven Explanation (RDE)
@@ 3.1.1. Causal Model
This diagram allows us to define target causal estimands such as $E[O\mid do(T)]$ (i.e., the deconfounded distribution after adjustment), which represents the expected model outcome under an intervention on $T$. In practice, causal models are often specified by experts, derived from controlled experiments, or inferred using causal discovery algorithms (Gong et al., 2024; Varando et al., 2024), which themselves rely on statistical associations in the observed data. We begin with a hypothesized causal model that assumes protected attributes act as confounders and explore different scenarios within this structure. Even when based on assumptions, such models provide a structured framework for reasoning under interventions, something that correlational methods often fail to do, especially in the presence of confounding (as illustrated by Simpson’s paradox (Pearl, 2022)). For instance, in Section 4.1, we consider the possibility that a loan applicant’s status influences the credit amount requested, an entirely plausible scenario in practice. The outcome $O$ may be either the model prediction $\hat{Y}$ or a derived quantity g(Y^)g($\hat${Y}). For example, in time-series settings, we define $O=|\hat{Y}-Y|$ as the residual error. This lets us ask: “Does a certain input perturbation systematically worsen performance for a particular group?”.
3.1.2. Rating metrics
@@
While RDE builds on prior rating methods, it repositions their function. Earlier works primarily used ratings to compare model performance across input conditions. RDE instead uses them as part of a structured explanation workflow: formulating a question (e.g., “Does treatment T affect outcome O?”), selecting a baseline, applying a metric, and interpreting the result as a potential explanation. This gives ratings an explanatory role, especially in helping stakeholders reason about model bias, instability, or robustness under different conditions. We use three metrics, each designed to answer a specific kind of question. The full implementation details are in Appendix A.
Weighted Rejection Score (WRS). Let $Z$ have $m$ groups and let $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ be a set of confidence levels (e.g., 95%,75%,60%), with corresponding weights $w_{k}$. For each unordered pair $(a,b)$ of groups and each confidence level $c_{k}$ compute the two-sample $t$-test statistic $t_{ab}^{(k)}$ and define the indicator
$I_{ab}^{(k)}=\begin{cases}1&\text{if }|t_{ab}^{(k)}|>t_{\mathrm{crit}}(c_{k},\,\nu_{ab}^{(k)})\\ 0&\text{otherwise,}\end{cases}$
where $t_{\mathrm{crit}}(c_{k},\nu)$ is the critical $t$-value at confidence $c_{k}$ with degrees of freedom $\nu_{ab}^{(k)}$. The WRS is the weighted sum of these rejections:
$\mathrm{WRS}\;=\;\sum_{k=1}^{K}w_{k}\sum_{1\leq a<b\leq m}I_{ab}^{(k)}.$
In our experiments we use $\mathcal{C}=\{0.95,0.75,0.60\}$ with weights $w=\{1.0,0.8,0.6\}$; the inner sum runs over all $\binom{m}{2}$ group pairs.
Average Treatment Effect (ATE). When the question is about the effect of a specific change (e.g., halving the requested loan amount), we estimate the ATE,
$\text{ATE}\;=\;\mathbb{E}[O\mid\text{do}(T=1)]\;-\;\mathbb{E}[O\mid\text{do}(T=0)],$
which is the average change in the model outcome under the intervention. In practice we estimate this with propensity-score matching (PSM) or G-computation when the treatment is continuous. Small ATE means the model is effectively insensitive to that change; a large ATE means the change shifts predictions or errors.
Deconfounded Impact Estimation (DIE%). An observed ATE can be misleading if a protected attribute $Z$ affects both $T$ and $O$. To quantify that distortion we compute
$\text{DIE\%}\;=\;100\times\frac{\lvert\text{ATE}_{\text{unadj}}-\text{ATE}_{\text{adj}}\rvert}{\lvert\text{ATE}_{\text{adj}}\rvert}$,
where $\text{ATE}_{\text{adj}}$ is the ATE after adjusting for $Z$ (via PSM or G-computation) and $\text{ATE}_{\text{unadj}}$ is the ATE before adjusting. If DIE% is large, much of the apparent effect of $T$ is actually explained by group differences; if DIE% is small, the treatment effect is robust to confounding.
These three measures play distinct roles in explanation: WRS measures distributional disparities across groups; ATE quantifies how much an intervention changes model outputs on average; DIE% quantifies the effect of confounders on the relationship between treatment and the outcome. We selected these metrics because they are straightforward to explain to non-technical stakeholders (regulators, applicants) and because together they let us rapidly triage whether a model’s behavior warrants a deeper, instance-level analysis (for example, using SHAP or counterfactuals).
3.1.3. RDE Workflow